OpenStudy (anonymous):

write y=0.81+25 in standard form

OpenStudy (solomonzelman):

Can you add 0.81+25?

OpenStudy (anonymous):

I don"t think so!

OpenStudy (solomonzelman):

Well.....

OpenStudy (solomonzelman):

Are you saying you don;t think so that you should add, or that you can add?

OpenStudy (anonymous):

I 'am not for sure exactly what to do sorry!

OpenStudy (solomonzelman):

Why not add 0.81+25 and set it equal to y

OpenStudy (anonymous):

I'am very weak in math! the choices are: -0.81x+y=25,0.81x+y=25,81x-100y=2500,and 0.81x-y=25.

OpenStudy (anonymous):

double check you copied the problem correctly according to the answer choices there should be an x in the original question

OpenStudy (anonymous):

write y=0.81x+25 in standard form

OpenStudy (anonymous):

sweet no standard form looks like this (Ax+By=C)...right now you have y on the correct side of the equation and B=1... you just have to get x to correct side

OpenStudy (anonymous):

you can subtract .81x from both sides you get .81x+y=25

OpenStudy (anonymous):

sorry -.81x+y=25

OpenStudy (anonymous):

thanks so much!!!!!!!!!!!!!!!!!!!!!!!!:)